OpenStudy (anonymous):

Select the augmented matrix for the following system of equations. y+2=3x 1/4z=z-1 z-8=x/2

OpenStudy (zzr0ck3r):

get them in the form ax+by=c

OpenStudy (zzr0ck3r):

so the first equation is -3x+y=2

OpenStudy (zzr0ck3r):

and then this will be one of the rows of the matrix -3 1 2

OpenStudy (anonymous):

What about the backwards equation? c=by+ax?

OpenStudy (zzr0ck3r):

o sorry I didn't see the z one of your rows will look like -3 1 0 2

OpenStudy (zzr0ck3r):

ok note we have x y z as rariables

OpenStudy (anonymous):

Does that require a separation between the 0 and 2 for each row?

OpenStudy (zzr0ck3r):

so the first equation can be written as -3x+1*y+0*z=2 does this make sense?

OpenStudy (anonymous):

Yes, that does.

OpenStudy (zzr0ck3r):

the second equation is 1/4z=z-1 can be written like (1/4)z-z=-1 and so -(3/4)z = -1 so we can write it like this 0*x+0*y-(3/4)z = -1 right?

OpenStudy (anonymous):